[Hallc_running] [New Logentry] RC Daily Update
sjjooste at jlab.org
sjjooste at jlab.org
Mon Dec 9 15:25:01 EST 2019
We finished the initial target install on Saturday afternoon (with help from the techs) and told MCC we were ready for beam permit by Saturday afternoon. After that there was a problem configuring the injector to deliver beam to Hall C, which was resolved by the OWL shift. It took OWL through DAY to slowly establish beam.
There were some issues with the 3H07A and 3H07B harps which were fixed by the expert (and Eric Pooser). Additionally, we discovered that the entire card with the target FSDs was masked, which meant we ran much of the startup without target-related FSD. Luckily, no major disasters happened, and it is unclear how this was possible in the first place...
After that we did the Bull's Eye scan, beam centering with the 2-hole target and ion chamber calibrations during SWING. With the beam checkout out of the way, we took our experiment-critical 2-pass optics data during OWL. We also attempted a BCM calibration but due to confusion about the operational restrictions during this procedure we decided to punt to conserve time and maximize the optics data we could take.
We took optics data until 5:30am, when a vacuum issue with the north linac stopped beam delivery. This issue is still ongoing. We did successfully collect all our 2-pass optics data we needed.
The original plan was to start preparing for 5-pass this morning while Dave Gaskell prepares to Moller to participate in the Spin Dance on Tuesday. We took the down as an opportunity to go into controlled access and do some necessary maintenance work:
* Prepare raster for 5-pass
* Setup HMS magnets
* Fix leak in reference cell
* Work on the polarimetry checkout
During this maintenance work, we noticed damage to the aluminum foil covering the Be window. We are working with safety to mitigate the issue and expect to fix this early SWING.
Assuming the accelerator is back up at this time, we can start preparing for 5pass for the spin dance.
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3750402
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running